### v1

@@ Abstract
Motivation: Medicinal synergy prediction is a powerful tool in drug discovery and development that harnesses the principles of combination therapy to enhance therapeutic outcomes by improving efficacy, reducing toxicity, and preventing drug resistance. While a myriad of computational methods has emerged for predicting synergistic drug combinations, a large portion of them may overlook the intricate, yet critical relationships between various entities in drug interaction networks, such as drugs, cell lines, and diseases. These relationships are complex and multidimensional, requiring sophisticated modeling to capture nuanced interplay that can significantly influence therapeutic efficacy.
Results: We introduce a salient deep hypergraph learning method, namely, Heterogeneous Entity Representation for MEdicinal Synergy prediction (HERMES), to predict anti-cancer drug synergy. HERMES integrates heterogeneous data sources, encompassing drug, cell line, and disease information, to provide a comprehensive understanding of the interactions involved. By leveraging advanced hypergraph neural networks with gated residual mechanisms, HERMES can effectively learn complex relationships/interactions within the data. Our results show HERMES demonstrates state-of-the-art performance, particularly in forecasting new drug combinations, significantly surpassing previous methods. This advancement underscores the potential of HERMES to facilitate more effective and precise drug combination predictions, thereby enhancing the development of novel therapeutic strategies.
@@ 1 Introduction
Exploring drug combinations leads a promising avenue for enhancing cancer treatment efficacy while minimizing toxicity and adverse reactions in modern medicine (Jia et al.,, 2009; Csermely et al.,, 2013). Combination therapies, involving multiple drugs administered as a single treatment regimen, offer potential benefits over traditional single-drug approaches, particularly under cancer and tumor treatment contexts (Chou,, 2006; O’Neil et al.,, 2016). Not only do they hold the promise of greater therapeutic efficacy, but present an opportunity to mitigate host toxicity and unwanted side effects, as the doses of drug combinations are often sub-mutagenic compared to individual drug doses. However, optimizing drug combinations can be challenging, as poorly chosen combinations may lead to adverse effects and sub-optimal outcomes (Hecht et al.,, 2009; Tol et al.,, 2009). Thus, there is a critical need to identify precise synergistic drug pairs tailored to different cancer types.
@@
Recently, numerous computational methods for predicting drug synergy have been proposed. Pioneering methods, including DeepSynergy (Preuer et al.,, 2018) and Matchmaker (Kuru et al.,, 2021), utilize deep neural networks with both the chemical properties of drugs and the gene expression profiles of cell lines. The deep tensor factorization model (Sun et al.,, 2020) combines tensor decomposition with neural networks to forecast the synergistic effects of drug combinations. Additionally, TransSynergy (Liu and Xie,, 2021) adopts a transformer network model using drug-target and protein-protein interaction data. DeepDDS (Wang et al.,, 2022) employs graph convolutional networks (GCNs) (Wu et al.,, 2020) and multi-layer perceptrons (MLP) for synergy prediction. The current state-of-the-art method HypergraphSynergy (Liu et al.,, 2022) made strides in this direction by incorporating hypergraph neural networks (HGNNs) (Feng et al.,, 2019; Bai et al.,, 2021) to model these dynamics in a more interconnected and multifaceted manner. Hypergraphs generalize graphs by allowing hyperedges to connect more than two nodes (Chen and Rajapakse,, 2020; Chen et al.,, 2021; Pickard et al.,, 2023), encode multidimensional (or high-order) correlations and connections. However, HypergraphSynergy has notable limitations in fully capturing the intricate nature of these interactions. More importantly, its ability to generalize, especially in the context of novel drug combinations, remains a challenge.
@@
In this article, we propose a novel deep hypergraph learning method – Heterogeneous Entity Representation for MEdicinal Synergy prediction (HERMES) – to enhance drug synergy prediction. HERMES distinguishes itself with its innovative strategy for integrating a variety of data sources, including drug chemical properties, cell line gene expressions, and interactions between drugs and indications. This integration is achieved through a heterogeneous hypergraph structure, which aids in assimilating extensive prior knowledge and enhances the model’s ability to generalize in novel contexts. Another notable breakthrough of HERMES is addressing the widespread issue of over-smoothing in message-passing networks (Li et al.,, 2018) by implementing a gated residual mechanism, which not only retains more information but also notably enhances the expressiveness of the network. Our empirical results highlight HERMES’s effectiveness, especially in novel scenarios, significantly surpassing HypergraphSynergy and establishing it as a leading solution. The key contributions of this article are (1) integration of varied knowledge sources (including drugs, cell lines, and diseases) into a scalable and heterogeneous model architecture; (2) enhancement of message passing over hypergraphs with gated residual mechanisms for augmented network expressiveness; (3) empirical evidence showcasing superior performance compared to previous methods, with a particular emphasis on robust generalization in novel contexts.
This article is structured into four sections. The main architecture of HERMES is detailed in Section 2. We assess the performance of HERMES (along with other representative drug synergy prediction methods) including a comprehensive ablation study in Section 3. Finally, we conclude by discussing future research directions in Section 4.
@@ 2.1 Overview of HERMES
Our framework conceptualizes drug synergy prediction under a hypergraph framework, treating it as a hyperedge prediction problem, where drugs, cell lines, and diseases are represented by nodes while synergistic drug–drug–cell line triplets and drug-disease pairwise relations are represented by hyperedges. Hyperedge prediction is a generalization of edge prediction on graphs (Chen and Liu,, 2023; Chen et al.,, 2023). This innovative approach unfolds through three interconnected phases: initialization, refinement, and consolidation (Figure 1). Each phase plays a pivotal role in processing and integrating varied data types, thereby synthesizing comprehensive information crucial for accurate drug synergy predictions. Drawing inspiration from HypergraphSynergy, our model introduces novel methodologies and integrates cutting-edge techniques to significantly enhance prediction accuracy and reliability at each phase of analysis.
@@ 2.2 Feature Initialization
The initialization phase is crucial in our synergy prediction methodology, as it involves acquisition of initial features for drugs, cell lines, and diseases. This phase is pivotal in not only gathering diverse data but ensuring these features are transformed to a uniform dimensionality, which is fundamental for the subsequent analyses in our hypergraph model.
Drug features. We generate drug features by utilizing molecular graphs, which are built from the Simplified Molecular Input Line Entry System (SMILES) representations of drugs (Kim et al.,, 2019). By employing the ConvMolFeaturizer method from the DeepChem library (Duvenaud et al.,, 2015), we construct a molecular graph for each drug, with atoms represented as nodes and bonds as edges. Subsequently, we employ advanced graph transformer networks (GTNs) on molecular graphs to enhance the representations of atoms and drugs. Denote the feature vector of atom $i$ as $\textbf{a}_{i}$. The updated feature (denoted as $\textbf{a}^{\prime}_{i}$) through a GTN is then computed as
@@
where $d$ is the latent size of each head, and $\textbf{W}_{3}$ and $\textbf{W}_{4}$ are learnable weight matrices. GTNs can effectively capture and model complex relationships and interactions within graph-structured data by assigning attention scores to neighboring nodes, enabling a more targeted aggregation of information compared to traditional convolutional methods such as GCNs (Shi et al.,, 2020). Finally, we compute the feature vector of a drug by aggregating all its updated atom features using maximum pooling. The choice of maximum pooling is motivated by the fact that drug molecular graphs may vary in the number of atom nodes. The resulting feature matrix for all drugs is denoted as $\textbf{X}_{\text{drug}}$.
@@ 2.3 Feature Refinement
In the refinement phase, we focus on transforming the obtained drug, cell line, and disease features into contextually enriched representations. This process is achieved through the construction of a dual-relation hypergraph and the implementation of an advanced feature enhancement technique, namely HGNNs with gated residual connections. These elements collectively elevate our model’s ability to identify complex interrelationships, setting the stage for more precise synergy predictions.
Hypergraph construction. We construct a novel dual-relationship hypergraph involving relationships between drugs, cell lines, and diseases. The hypergraph contains two types of hyperedges: (1) drug-drug-cell line triplets (third-order interactions), which capture the potential synergistic effects between specific drugs and cell lines; (2) drug-disease relationship (pairwise interactions or traditional edges), which represents associations between drugs and their indications (medical conditions they treat), providing insights into the therapeutic spectrum of drugs. This dual-relationship hypergraph, which has never been considered in previous methods, is pivotal in enriching the feature representations by integrating the various features of drugs, cell lines, and diseases.
@@
where D and E are diagonal matrices of node and hyperedge degrees, respectively. H is the incidence matrix of the hypergraph, containing binary values to indicate the presence or absence of any node in any hyperedge, and $\textbf{W}_{6}$ is learnable weight matrix. Here, the superscripts ${-1}$ and $\top$ denote matrix inversion and transposition, respectively. To enhance the expressive capacity of our network and extract more intricate patterns from the hypergraph, we increase the number of layers in the HGNN. However, this approach leads to a well-known challenge in graph neural networks, i.e., over-smoothing (Cai and Wang,, 2020).
@@ 2.5 HERMES and HypergraphSynergy
Below, we show that HERMES significantly outperforms HypergraphSynergy in two drug synergy datasets, especially under the context of predicting new drug combinations.
@@ 3.1 Datasets
  • Gene expression in cancer cell lines: Data on gene expression in cancer cell lines are sourced from the Cell Lines Project within the COSMIC database (Forbes et al.,, 2015). In this context, we specifically considered the expression values of $651$ genes related to the COSMIC cancer gene census. These expression values are subjected to logarithmic (log2) transformation and z-score normalization.
  • Indications for drug therapy in diseases: The indication dataset is sourced from PrimeKG (Chandak et al.,, 2023), a multi-modal knowledge graph designed to support precision medicine. PrimeKG integrates data from $20$ high-quality biomedical resources, including DisGeNET and DrugBank. It provides detailed relationships between drugs and diseases, covering indications, contraindications, and off-label uses. This dataset captures how drugs target disease-associated molecular perturbations, facilitating analyses in drug repurposing and therapeutic strategies
We performed a comprehensive data preprocessing on two primary drug synergy datasets – NCI-ALMANAC and O’Neil. To ensure the quality and relevance of the data, we excluded cell lines lacking gene expression information and drugs without SMILES details. Following this preprocessing phase, the NCI-ALMANAC dataset comprises $74,139$ measurement samples of ComboScores for $87$ drugs across $55$ cancer cell lines, while the O’Neil dataset encompasses $18,950$ samples of Loewe synergy scores for $38$ drugs and $39$ cancer cell lines (Table 1). Subsequently, we removed drugs from the ‘drug-disease dataset’ that are not present in the aforementioned drug synergy datasets. This led to the extraction of indications for $111$ diseases corresponding to $111$ drugs within the NCI-ALMANAC dataset and $222$ diseases corresponding to $222$ drugs within the O’Neil dataset.
Table 1: Statistics of the two datasets.
Dataset #Drugs #Targets #Samples
NCI-ALMANAC 87 55 74139
O’Neil 38 39 18950
@@ 3.3 Experiment Setup
In this study, each dataset is initially divided into two distinct sets: a training set, accounting for $90\%$ of the total data, and a test set comprising the remaining $10\%$. The training set undergoes a rigorous five-fold cross-validation process. This process is structured in three unique partitioning strategies to ensure comprehensive evaluation:
@@
  • Target: Samples are stratified by target cell line, ensuring each fold’s validation set contains unique cell lines not present in its training set.
@@
Additionally, for the classification task, synergy scores are converted to binary outcomes. Following established protocols (Preuer et al.,, 2018; Sun et al.,, 2020), a threshold of 30 is used. Scores above this threshold indicate a positive synergy, while scores below are deemed negative. Ultimately, the model’s performance is evaluated using key metrics: area under the receiver operating characteristic curve (AUROC) and area under the precision-recall curve (AUPRC). Both metrics provide comprehensive view of the model’s effectiveness in precision and recall aspects.

### v2

@@ Abstract
Motivation: Forecasting the synergistic effects of drug combinations facilitates drug discovery and development, especially regarding cancer therapeutics. While numerous computational methods have emerged, most of them fall short in fully modeling the relationships among clinical entities including drugs, cell lines, and diseases, which hampers their ability to generalize to drug combinations involving unseen drugs. These relationships are complex and multidimensional, requiring sophisticated modeling to capture nuanced interplay that can significantly influence therapeutic efficacy.
Results: We present a novel deep hypergraph learning method named Heterogeneous Entity Representation for MEdicinal Synergy prediction (HERMES) to predict the synergistic effects of anti-cancer drugs. Heterogeneous data sources, including drug chemical structures, gene expression profiles, and disease clinical semantics, are integrated into hypergraph neural networks equipped with a gated residual mechanism to enhance high-order relationship modeling. HERMES demonstrates state-of-the-art performance on two benchmark datasets, significantly outperforming existing methods in predicting the synergistic effects of drug combinations, particularly in cases involving unseen drugs.
@@ 1 Introduction
Exploring drug combinations leads a promising avenue for enhancing cancer treatment efficacy while minimizing toxicity and adverse reactions in modern medicine (Jia et al.,, 2009; Csermely et al.,, 2013; Wang et al.,, 2012; Foucquier and Guedj,, 2015). Combination therapies, involving multiple drugs administered as a single treatment regimen, offer potential benefits over traditional single-drug approaches, particularly under cancer and tumor treatment contexts (Chou,, 2006; O’Neil et al.,, 2016). Not only do they hold the promise of greater therapeutic efficacy, but present an opportunity to mitigate host toxicity and unwanted side effects, as the doses of drug combinations are often sub-mutagenic compared to individual drug doses. However, optimizing drug combinations can be challenging, as poorly chosen combinations may lead to adverse effects and sub-optimal outcomes (Hecht et al.,, 2009; Tol et al.,, 2009). Thus, there is a critical need to identify precise synergistic drug pairs tailored to different cancer types.
@@
Recently, numerous computational methods for predicting drug synergy have been proposed. Pioneering methods, including DeepSynergy (Preuer et al.,, 2018) and Matchmaker (Kuru et al.,, 2021), utilize deep neural networks with both the chemical properties of drugs and the gene expression profiles of cell lines. The deep tensor factorization model (Sun et al.,, 2020) combines tensor decomposition with neural networks to forecast the synergistic effects of drug combinations. Additionally, TransSynergy (Liu and Xie,, 2021) adopts a transformer network model using drug-target and protein-protein interaction data. DeepDDS (Wang et al.,, 2022) employs graph convolutional networks (GCNs) (Wu et al.,, 2020) and multi-layer perceptrons (MLP) for synergy prediction. The current state-of-the-art method HypergraphSynergy (Liu et al.,, 2022) made strides in this direction by incorporating hypergraph neural networks (HGNNs) (Feng et al.,, 2019; Bai et al.,, 2021) to model these dynamics in a more interconnected and multifaceted manner. Hypergraphs generalize graphs by allowing hyperedges to connect more than two nodes (Chen and Rajapakse,, 2020; Chen et al.,, 2021; Pickard et al.,, 2023), encode multidimensional (or high-order) correlations and connections. However, these existing methods often fall short in effectively predicting drug synergy due to their neglect of higher-order drug interactions or lack of important biomedical knowledge. DeepDDS focuses solely on pairwise drug interactions using GCNs, and HypergraphSynergy, although it considers higher-order drug-drug-cell line combinations, fails to incorporate additional biomedical interactions, which hinders its generalization, particularly involving unseen drug molecules.
@@
In this article, we propose a novel deep hypergraph learning method – Heterogeneous Entity Representation for MEdicinal Synergy prediction (HERMES) – to enhance drug synergy prediction. HERMES distinguishes itself with its innovative strategy for integrating a variety of data sources, including drug chemical properties, cell line gene expressions, and interactions between drugs and indications. This integration is achieved through a heterogeneous hypergraph structure, which aids in assimilating extensive prior knowledge and enhances the model’s ability to generalize in novel contexts. Incorporating information about drug indications is particularly important, as some drugs may function through similar molecular mechanisms in different diseases (e.g., Bevacizumab is effective against various cancers, including colorectal cancer and non-small cell lung cancer). In contrast, drugs like glucocorticoids demonstrate different mechanisms across diseases due to distinct biological pathways, such as their anti-inflammatory action in rheumatoid arthritis versus their apoptosis-inducing role in leukemia. This approach allows for more comprehensive capturing of the varied biological contexts in which drug combinations achieve synergetic effects. Another notable breakthrough of HERMES is addressing the widespread issue of over-smoothing in message-passing networks by implementing a gated residual mechanism (Li et al.,, 2018), which not only retains more information but also notably enhances the expressiveness of the network.
Our empirical results highlight HERMES’s effectiveness, especially in novel scenarios, surpassing HypergraphSynergy and establishing it as a leading solution. The key contributions of this article are (1) integration of varied knowledge sources (including drugs, cancer cell lines, and disease indications) into a scalable and heterogeneous model architecture; (2) enhancement of message passing over hypergraphs with gated residual mechanisms for augmented network expressiveness; (3) empirical evidence showcasing superior performance compared to previous methods, highlighting robust generalization in novel contexts, with an average improvement of 5% in evaluation metrics over the state-of-the-art methods. This article is structured into four sections. The main architecture of HERMES is detailed in Section 2. We assess the performance of HERMES (along with other representative drug synergy prediction methods) including a comprehensive ablation study in Section 3. Finally, we conclude by discussing future research directions in Section 4.
@@ 2.1 Overview of HERMES
Our framework conceptualizes drug synergy prediction under a hypergraph framework, treating it as a hyperedge prediction problem, where drugs, cell lines, and diseases are represented by nodes while synergistic drug–drug–cell line triplets and drug-disease pairwise relations are represented by hyperedges. Hyperedge prediction is a generalization of edge prediction on graphs (Chen and Liu,, 2023; Chen et al.,, 2023; Zhang et al.,, 2018; Sharma et al.,, 2021; Yadati et al.,, 2020; Kumar et al.,, 2020). This innovative approach unfolds through three interconnected phases: initialization, refinement, and consolidation (Figure 1). Each phase plays a pivotal role in processing and integrating varied data types, thereby synthesizing comprehensive information crucial for accurate drug synergy predictions. Drawing inspiration from HypergraphSynergy, our model introduces novel methodologies and integrates cutting-edge techniques to significantly enhance prediction accuracy and reliability at each phase of analysis.
@@ 2.2 Feature Initialization
The initialization phase is crucial in our synergy prediction methodology, as it involves the acquisition of initial features for drugs, cell lines, and diseases. This phase performs modality-specific representation learning to transform the raw features of different modalities into a common feature space. This alignment process is critical for the subsequent hypergraph learning as it ensures consistency in the representations, allowing for effective integration of features across modalities.
Drug features. We generate drug features by utilizing molecular graphs, which are built from the Simplified Molecular Input Line Entry System (SMILES) representations of drugs (Kim et al.,, 2019). By employing the ConvMolFeaturizer method from the DeepChem library (Duvenaud et al.,, 2015), we construct a molecular graph for each drug, with atoms represented as nodes and bonds as edges. Subsequently, we employ advanced graph transformer networks (GTNs) (Yun et al.,, 2019) on molecular graphs to enhance the representations of atoms and drugs. Denote the feature vector of atom $i$ as $\textbf{a}_{i}$. The updated feature (denoted as $\textbf{a}^{\prime}_{i}$) through a GTN is then computed as
@@
where $d$ is the latent size of each head, and $\textbf{W}_{3}$ and $\textbf{W}_{4}$ are learnable weight matrices. GTNs can flexibly capture the intricate and long-range dependencies among drugs’ atoms by adaptively learning attention weights among them, offering more comprehensive drug representations over traditional GCNs which are often limited by local receptive fields. Finally, we compute the feature vector of a drug by aggregating all its updated atom features using maximum pooling. The choice of maximum pooling is motivated by the fact that drug molecular graphs may vary in the number of atom nodes. The resulting feature matrix for all drugs is denoted as $\textbf{X}_{\text{drug}}$.
@@ 2.3 Feature Refinement
In the refinement phase, we focus on transforming the obtained drug, cell line, and disease features into contextually enriched representations. This process is achieved through the construction of a dual-relation hypergraph and the implementation of an advanced feature enhancement technique, namely HGNNs with gated residual connections (Li et al.,, 2018). These elements collectively elevate our model’s ability to identify complex interrelationships, setting the stage for more precise synergy predictions.
Hypergraph construction. We construct a novel dual-relationship hypergraph involving relationships between drugs, cell lines, and diseases. The hypergraph contains two types of hyperedges: (1) drug-drug-cell line triplets (third-order interactions), which capture the potential synergistic effects between specific drugs and cell lines; (2) drug-disease relationships (pairwise interactions or traditional edges), which represent associations between drugs and their corresponding disease indications. In this context, diseases refer specifically to drug indications rather than the broader cancer types represented by cell lines. Therefore, we do not explicitly construct hyperedges between cell lines and diseases. This dual-relationship hypergraph, which has never been considered in previous methods, is pivotal in enriching the feature representations by integrating the various features of drugs, cell lines, and diseases.
@@
where D and E are diagonal matrices of node and hyperedge degrees, respectively. H is the incidence matrix of the hypergraph, containing binary values to indicate the presence or absence of any node in any hyperedge, and $\textbf{W}_{6}$ is the learnable weight matrix. Here, the superscripts ${-1}$ and $\top$ denote matrix inversion and transposition, respectively. To enhance the expressive capacity of our network and extract more intricate patterns from the hypergraph, we increase the number of layers in the HGNN. However, this approach leads to a well-known challenge in graph neural networks, i.e., over-smoothing (Chen et al.,, 2020).
@@ 2.5 HERMES and HypergraphSynergy
To mitigate the risk of overfitting, HERMES also includes various regularization techniques, including dropout layers, weight decay, and early stopping mechanism. Below, we show that HERMES significantly outperforms HypergraphSynergy in two drug synergy datasets, especially under the context of predicting new drug combinations.
@@ 3.1 Datasets
  • Gene expression in cancer cell lines: Data on gene expression in cancer cell lines are sourced from the Cell Lines Project within the COSMIC database (Forbes et al.,, 2015). In this context, we specifically considered the expression values of $651$ genes related to the COSMIC cancer gene census. These expression values are subjected to logarithmic (log2) transformation and z-score normalization. We also utilized the CCLE (Barretina et al.,, 2012) and GDSC (Yang et al.,, 2012) databases as alternative data sources (see details in Supplementary Data).
  • Drug indications: The drug-indication annotations are sourced from PrimeKG (Chandak et al.,, 2023), a multimodal knowledge graph designed for precision medicine. PrimeKG integrates information from various high-quality biomedical resources, such as DisGeNET Piñero et al., (2020) and DrugBank Wishart et al., (2018), as well as data from different biological scales, including disease pathways and phenotypes. It provides detailed relationships between drugs and diseases, covering indications, contraindications, and off-label uses. This comprehensive knowledge base is particularly valuable for identifying potential new applications for drugs and optimizing treatment approaches based on the underlying molecular mechanisms of diseases.
We performed a comprehensive data preprocessing on two primary drug synergy datasets – NCI-ALMANAC and O’Neil. To ensure the quality and relevance of the data, we excluded cell lines lacking gene expression information and drugs without SMILES details. Following this preprocessing phase, the NCI-ALMANAC dataset comprises $74,139$ measurement samples of ComboScores for $87$ drugs across $55$ cancer cell lines, while the O’Neil dataset encompasses $18,950$ samples of Loewe synergy scores for $38$ drugs and $39$ cancer cell lines (Table 1). Subsequently, we removed drugs from the ‘drug-disease dataset’ that are not present in the aforementioned drug synergy datasets. This led to the extraction of indications for $82$ diseases corresponding to $37$ drugs within the NCI-ALMANAC dataset and $42$ diseases corresponding to $12$ drugs within the O’Neil dataset.
@@ 3.2 Baselines
  • NHP (Yadati et al.,, 2020): It is a GCN-based model specifically designed for hypergraphs to capture complex, higher-order relationships among multiple nodes. NHP employs hyperlink-aware GCN layers to transform hyperedges into clique expansions, which enables the modeling of multi-way interactions among drugs and cell lines in the drug synergy prediction task.
@@ 3.3 Experiment Setup
In this study, we trained and validated the models independently using the NCI-ALMANAC and O’Neil datasets. For each dataset, it is initially divided into two distinct sets: a training set, accounting for $90\%$ of the total data, and a test set comprising the remaining $10\%$. The training set undergoes a rigorous five-fold cross-validation process. This process is structured in three unique partitioning strategies to ensure comprehensive evaluation:
@@
  • CLine: Samples are stratified by target cell line, ensuring each fold’s validation set contains unique cell lines not present in its training set.
@@
Table 1: Statistics of the two datasets.
Dataset #Drugs #CLine #Samples
NCI-ALMANAC 87 55 74139
O’Neil 38 39 18950
Additionally, for the classification task, synergy scores are converted to binary outcomes. Following established protocols (Preuer et al.,, 2018; Sun et al.,, 2020), a threshold of 30 is used. Scores above this threshold indicate a positive synergy, while scores below are deemed negative. Ultimately, the model’s performance is evaluated using key metrics: area under the receiver operating characteristic curve (AUROC), area under the precision-recall curve (AUPRC), and F1-score. These three metrics together provide a comprehensive view of the model’s effectiveness in terms of precision, recall, and overall balance between precision and recall.